A car is traveling 30 mph faster than a bus, in the time that it takes the bus to travel 105 miles the car travels 195 miles, what is the speed of the car?
Im having trouble ive tried utilizing the v=d/t equation but i keep getting confused in my chart. Your help would be appreciated.

Answer by mananth(16946)

A car is traveling 30 mph faster than a bus, in the time that it takes the bus to travel 105 miles the car travels 195 miles, what is the speed of the car?

bus speed be x mph
bus distance = 105
bus time = 105/x
car speed = x+30 mph
car distance = 195
car time = 195/(x+30)
..
car time = bus time
195/(x+30)= 105/x
195x=105(x+30)
195x=105x+3150
195x-105x= 3150
90x = 3150
/90
x= 35 mph speed of bus
speed of car = 30+35 = 65 mph
...
m.ananth@hotmail.ca